Find the coefficient of variation, k. Write a Function that models each variation.
x = 5 and y = 3 when z = 7. z varies directly with x and inversely with y.

Answer by stanbon(75887)

Find the coefficient of variation, k. Write a Function that models each variation.
x = 5 and y = 3 when z = 7. z varies directly with x and inversely with y.
-----
z = k*x/y
-------------------
7 = k*5/3
k = 21/5
k = 4.2
-----------
Cheers,
Stan H.
--------------